OpenStudy (luckycoins888):

fan = medal The regular price of an item is $84. The item is on sale with a discount of $12. What is the discount rate? A. 0.143% B. 1.43% C. 14.3% D. 143%

OpenStudy (theedwardsfamily):

i think C for sure

OpenStudy (theedwardsfamily):

@luckycoins888

OpenStudy (anonymous):

\[ \frac{$12}{$84}=\frac{x\%}{100\%} \]

OpenStudy (anonymous):

This will tell you what percent of the original price it was.

OpenStudy (anonymous):

If we are talking about something like \(y\%\) off, then you would use: \[ $84\times (100-y)\%=$12 \]

OpenStudy (anonymous):

Basically: \[ \frac{$12}{$84} =\frac{(100-y)\%}{100\%} \]

OpenStudy (anonymous):

*he he he*

OpenStudy (anonymous):

@omg

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

got u

OpenStudy (anonymous):

@crazychickengirl123 You're really crazy